Plane sound wave of frequency 500 Hz fall normally on a smooth wall.At what distance from the wall will the air particles have maximum and minimum amplitude of vibration and the velocity of sound in air may be taken as 340 meter per second

Grade:12

1 Answers

Arun
4 years ago
 
Shortest distance should be lambda/4
 
Now
Lambda = v/f = 330/500 = 0.66
 
 
Hence distance = 0.66/4 = 0.165 m
 
Hope it helps